Β A demographer estimates that the population of inmates at the Nsawam prisons will increase at the rate of 298𝑒0.8π‘₯ per year in x years. Currently there are about 7800 inmates at the prison. How many inmates should the prison authorities expect 10 years from now?


1
Expert's answer
2021-11-09T17:30:22-0500

"\frac{dN}{dx}=298e^{0.8x}\\
where \space N(x=0) = 7800\\
dN = 298e^{0.8x} dx\\
N=\frac{298e^{0.8x}}{0.8}+C\\
At \space x=0\\
7800= \frac{298}{0.8}+C \implies C = 7427.5\\
N=\frac{298e^{0.8x}}{0.8}+7427.5\\
N(10)=\frac{298e^{0.8*10}}{0.8}+7427.5\\
N= 1117834.35017\\
N=1117835 \space inmates"